OpenStudy (anonymous):

y=p-2qx^2 p=-10 q=2 x=-5

OpenStudy (anonymous):

@bibby is it -10-2(2+-5^2)

OpenStudy (bibby):

put the parentheses around the x^2

OpenStudy (bibby):

\[\huge -10-[2*2*(-5)^2]\]

OpenStudy (anonymous):

-90?

OpenStudy (bibby):

-10 - [4*25] -10 - 100 =-10 - 100 = -110

OpenStudy (anonymous):

so is it -110

OpenStudy (bibby):

it is

OpenStudy (anonymous):

thnx a million

OpenStudy (bibby):

no probs a billion

OpenStudy (anonymous):

LOL